Suppose the Japanese yen exchange rate is ¥77 = $1, and the British pound exchange rate is £1 = $1.61.

a.

What is the cross-rate in terms of yen per pound? (Round your answer to 2 decimal places, e.g., 32.16.)

  Cross-rate ¥/£    

  

b.

Suppose the cross-rate is ¥126 = £1. What is the arbitrage profit per dollar used?
Answer #1

a]

Cross-rate ¥/£ = (¥/$) * ($/£)

Cross-rate ¥/£ = 77 * 1.61 = ¥123.97 / £.

b]

Let us say we start with $1.

This is converted into £.  £ received = 1 / 1.61 =  £0.6211.

£0.6211 is converted into ¥. ¥ received = 0.6211 * 126 = ¥78.2609.

¥78.2609 is converted back into $. $ received =  78.2609 / 77 = $1.0164.

Arbitrage profit =  $1.0164 - $1 = $0.0164.